# The Polar Express: Optimal Matrix Sign Methods and Their Application to the Muon Algorithm

### v2

@@ 5.2   Training GPT-2
In our final experiment, we compare the performance of using our Polar Express method given in Algorithm 1 inside the Muon algorithm versus Jordan’s [22] and You’s [9] methods.
Our experimental setup is based on the modified nanogpt code of Jordan [21]. We train two different GPT-2 models:
[…]
and a vocabulary size of $50{,}257$, using a context length of $1024$. Training is performed on 1B tokens from the FineWeb dataset [2], using a batch size of 32 and a single epoch. All models are trained with mixed precision (bfloat16) on 4 H100 GPUs. For all methods we use the learning rate schedule proposed in [21], consisting of a constant phase for the first 40% of training steps followed by a linear decay. All methods for the matrix sign computations are performed in float16b precision and use five iterations.
We apply Muon selectively to certain layers of the model. Following the nano-gpt implementation [21], we assign Muon to all parameters with at least two dimensions (typically weight matrices, and excluding RMS norm parameters), excluding the embeddings, unembeddings, and the attention head. These excluded parameters are instead optimized with AdamW.
Figure 6: Training a GPT-2 (124M) model on 1 Billion tokens of the Fineweb data set [2]. The Legend muon-<name> refers to using muon with the <name> method for computing $\operatorname*{polar}(\bm{M})$. Top Left: The final validation loss vs. the learning rate. The final best validation losses for each method were, in reverse order, adamw: $4.197$, muon-Jordan: $3.639$, muon-You: $3.629$ and muon-PolarExp: $3.588$. Bottom Left: The final training loss vs the learning rate. Top Right: Validation loss vs. number of iterations. Bottom Left: validation loss vs. time, plotting each method with its best learning rate.
@@
The aim of this section is to prove Theorem 4.1. We begin with a result that provides a few essential properties for the the polynomial solving (7) when $T=1$. This result is known as Chebyshev’s theorem [10] or the equioscillation theorem [46, Chapter 10].
@@
By Chebyshev’s Theorem [1, 10, 12] it is sufficient to show that $\mathbb{P}_{d}^{\operatorname*{odd}}$ satisfies the Haar condition: any non-zero $p\in\mathbb{P}_{d}^{\operatorname*{odd}}=\mbox{span}\{x,\ldots,x^{3},\ldots,x^{2q+1}\}$ can have at most $q$ roots in $[\ell,u]$.
Since $\deg(p)=d=2q+1$ we know that $p$ can have at most $12q+$ 1 roots in $\mathbb{R}$. However, since $p(0)=0$ and $p(x)=-p(-x)$ we know that $p$ has one root at zero, and the remaining roots come in symmetric pairs $(x,-x)$. Because of this, $p$ can have at most $q$ roots in the positive orthant, and thus it can have at most $q$ roots in $[\ell,u]\subset(0,\infty)$. Hence, $\mathbb{P}_{d}^{\operatorname*{odd}}$ satisfies the Haar condition, which yields the desired result.
@@
Note that $e(x)$ has at most $2q$ stationary points in $\mathbb{R}$, since its derivative $e^{\prime}(x)$ is a polynomial of degree $2q$. Furthermore, since $p^{\star}$ is odd, we have that $e^{\prime}(x)=-p^{\prime}(x)$ is even of degree $2q$, and thus can have at most $q$ stationary points contained in $(0,+\infty)$. Hence, there can be at most $q$ stationary points of $e(x)$ inside the interval $[\ell,u]$.
By Lemma A.1 there are $q+2$ points $x_{0},\ldots,x_{q+1}\in[\ell,u]$ where $e(x)$ is maximized or minimized in $[\ell,u]$. These points are either stationary points or they are endpoints of the interval $[\ell,u]$. Let $n_{\text{ext}}$ be the number of stationary points and $n_{\text{stat}}$ be the number of endpoints in the set $\{x_{0},\ldots,x_{q+1}\}$. Since a point can be both a stationary point and an endpoint we have $q+2\leq n_{\text{end}}+n_{\text{stat}}$. However, $n_{\text{end}}\leq 2$ and $n_{\text{stat}}\leq q$, which follows from the previous paragraph where we showed that there are at most $q$ stationary points of $e(x)$ in $[\ell,u]$. So $n_{\text{end}}+n_{\text{stat}}\leq q+2$, and consequently we must have $n_{\text{end}}=2$ and $n_{\text{stat}}=q$, as required.
Step 2: $x=\ell$ is a maximum of $e(x)$ on the interval $[\ell,u]$
By Lemma A.1 and the discussion from Step 1, we know that $|e(x)|$ is maximized at $q+2$ points inside $[\ell,u]$ and $q$ of these points are contained inside the open interval $(\ell,u)$. Hence, $x=\ell$ must either be a maximum or a minimum of $e(x)$. We will show that $x=\ell$ must be a maximum by contradiction.
Suppose $x=\ell$ was a minimum of $e(x)$ on $[\ell,u]$. First note that $p^{\star}$ is trivially non-negative on $[\ell,u]$, or else $p(x)=0$ would be a better polynomial. Hence, since $p^{\star}(0)=0$ we must have ${p^{*}}^{\prime}(\delta)>0$ for some $\delta\in[0,\ell]$, or else the zero polynomial $p(x)=0$ would be a better approximation. Hence, for some $\delta\in[0,\ell]$ we have $e^{\prime}(\delta)<0$.
We must also have $e^{\prime}(\ell)\geq 0$ or else $x=\ell$ is not a minimum of $e(x)$. Since $e^{\prime}(\delta)<0$ for some $\delta\in[0,\ell]$ and $e^{\prime}(\ell)\geq 0$, by the intermediate value theorem there exists a point $x^{*}\in[0,\ell]$ such that $e^{\prime}(x^{*})=0$. However, by the discussion above we know that all stationary points of $e$ are contained inside the open interval $(\ell,u)$. Hence, $x=\ell$ cannot be a minimum of $e(x)$ on $[\ell,u]$. However, by Step 1 we know that the endpoints of $[\ell,u]$ must be either minima or maxima of $e(x)$. Hence, $x=\ell$ is a maximum of $e(x)$ on $[\ell,u]$.
@@
Since $e(x)$ has a maximum in $[\ell,u]$ at $x=\ell$, we have $p^{\star}(\ell)=\min\limits_{x\in[\ell,u]}p^{\star}(x)$. The other two equalities are immediate consequences of the equioscillation property of $p^{\star}$ Lemma A.1 and that $x=\ell$ is a minimum of $p^{\star}$ over the set $[\ell,u]$. ∎
@@
We now proceed with the proof of (11), which will be by induction. The proof for $T=1$ is an immediate consequence of Lemma A.2 and we also have $p^{\star}(\ell)=\ell_{2}$ by (10). Now suppose the result is true for all $t\leq T-1$. For $t=1,\ldots,T-1$, note that the image of $p_{t}$ on $[\ell_{t},u_{t}]$ is exactly $[\ell_{t+1},u_{t+1}]$ by i). Hence, if we define $g(x):=p_{T-1}\circ\cdots\circ p_{1}(x)$, then the image of $g$ on $[\ell,u]$ is $[\ell_{T},u_{T}]$. Furthermore, by i) we also have $g(\ell)=\ell_{T}$. Pick any $f$ such that $f\neq g$ and
@@
for some $\widetilde{p}_{1},\ldots,\widetilde{p}_{T-1}\in\mathbb{P}_{d}^{\operatorname*{odd}}$. Let the image of $f$ on $[\ell,u]$ be $[a,b]$. We will prove that $\frac{a}{b}\leq\frac{\ell_{T}}{u_{T}}$ by contradiction.
Suppose $\frac{a}{b}>\frac{\ell_{T}}{u_{T}}$. Define $c=\frac{2}{a+b}$. Then, the image of the scaled function $cf$ on $[\ell,u]$ is $[ca,cb]$ and $cf$ satisfies
@@
Then Algorithm 2 returns $\bm{X}_{T}=p^{\star}(\bm{M})$. Let $h\in\mathbb{P}_{q}$ be $[q/0]$ Padé-approximant to $(1-x)^{-1/2}$ [25, Section 3] and define $p(x)=xh(1-x^{2})\in\mathbb{P}_{d}^{\operatorname*{odd}}$. Define $f=p\circ\cdots\circ p$ as the composition of $p$ with itself $T$ times. Then, by Theorem 4.1 and [25, Theorem 3.1] we have
@@ Appendix D Remez algorithm
Theorem 4.1 shows that we can solve (7) by greedily choosing the optimal approximation $p_{t}\in\mathbb{P}_{d}^{\operatorname*{odd}}$ for each interval $[\ell_{t},u_{t}]$ for $t=1,\ldots,T$. In this section, we outline how the Remez algorithm [37, 38] can be used to approximate $p_{t}$.
@@
We begin with the case when $d=3$. In this case, there is a simple closed form for the optimal odd polynomial $p^{\star}\in\mathbb{P}_{3}^{\operatorname*{odd}}$; see [11]. On a given interval $[\ell,u]$, the optimal approximation to the constant function $x\mapsto 1$ is given by the scaled and shifted Newton-Schulz polynomial $p_{\operatorname*{NS}}(x)=\frac{3}{2}x-\frac{1}{2}x^{3}$:
@@
$p^{\star}(x)=\beta p_{\operatorname*{NS}}(\alpha x),\text{ where }\alpha=\sqrt{\frac{3}{u^{2}+lu+\ell^{2}}}\text{ and }\beta=\frac{4}{2+\ell u(\ell+u)\alpha^{3}}.$
One can verify that this polynomial satisfies the equioscillation condition from Lemma A.1 at $x=\ell,\frac{1}{\alpha},u$ and therefore necessarily has to be the optimal approximation from $\mathbb{P}_{3}^{\operatorname*{odd}}$. Unfortunately, for larger $d$ finding closed form expressions for optimal approximations from $\mathbb{P}_{d}^{\operatorname*{odd}}$ becomes challenging. In fact, to the best of our knowledge the optimal approximation is not known for $d\geq 5$. However, we can approximate the optimal polynomial using the Remez algorithm.
@@
Let $d=2q+1$. Recall that from Lemma A.1 the optimal polynomial must satisfy the equioscillation property $q+2$ times. The Remez algorithm is an iterative algorithm that finds the equioscillation points $A=\{x_{0},\ldots,x_{q+1}\}$ from Lemma A.1 by iteratively refining a sequence of trial points $A^{(k)}=\{x_{0}^{(k)},\ldots,x_{q+1}^{(k)}\}$ so that $A^{(k)}$ converges to $A$. From the sequence of trial points $A^{(k)}$ the algorithm also finds a sequence of polynomials $p^{(k)}$ so that $p^{(k)}$ converges to the optimal polynomial. The convergence is extremely fast, and usually 10 iterations is sufficient to converge to the optimal polynomial up to double precision machine epsilon [37]. More commonly, the Remez algorithm is used to find optimal polynomial approximations to general continuous functions where $d\approx 100$ or even $d\approx 1000$. However, because the polynomial we build to approximate $\operatorname*{sign}(x)$ is a composition of polynomials, each of which has a low degree, in our setting the degree $d$ is small, usually $d=5$. For $d=5$ the Remez algorithm simplifies significantly. We now describe this simplified algorithm.
@@
Recall that as stated in Lemma A.1, the unique optimal approximation $p^{\star}\in\mathbb{P}_{5}^{\operatorname*{odd}}$ satisfies the equioscillation property at four points. The Remez algorithm first starts with a trial set $A^{(1)}=\{x_{0}^{(1)},x_{1}^{(1)},x_{2}^{(1)},x_{3}^{(1)}\}\subset[\ell,u]$ which ideally should come close to satisfying the equioscillation property. Since we know that $\ell$ and $u$ must be equioscillation points we always set $x_{0}^{(k)}=\ell$ and $x_{3}^{(k)}=u$ for all $k$. $x_{2}^{(1)}$ and $x_{3}^{(1)}$ are chosen to be $\frac{1}{4}\ell+\frac{3}{4}u$ and $\frac{3}{4}\ell+\frac{1}{4}u$, since we observe that as $\ell\to u$ these are approximately the other equioscillation points. Next, the algorithm solves the following system of equations
$a_{1}x_{i}^{(1)}+b_{1}(x_{i}^{(1)})^{3}+c_{1}(x_{i}^{(1)})^{5}+(-1)^{i}E_{1}=\operatorname*{sign}(x)=1,\quad i=0,1,2,3,$ (17)
for the unknowns $a_{1},b_{1},c_{1},$ and $E_{1}$. Recalling that $x_{0}^{(1)}=\ell$ and $x_{3}^{(1)}=u$, (17) can equivalently be written as a system of linear equations
$\begin{bmatrix}\ell&\ell^{3}&\ell^{5}&1\\ x_{1}^{(1)}&(x_{1}^{(1)})^{3}&(x_{1}^{(1)})^{5}&-1\\ x_{2}^{(1)}&(x_{2}^{(1)})^{3}&(x_{2}^{(1)})^{5}&1\\ u&u^{3}&u^{5}&-1\end{bmatrix}\begin{bmatrix}a\\ b\\ c\\ E\end{bmatrix}=\begin{bmatrix}1\\ 1\\ 1\\ 1\end{bmatrix}.$ (18)
Once we have solved for $a_{1},b_{1},c_{1},$ and $E_{1}$ we set $p_{1}(x)=a_{1}x+b_{1}x^{3}+b_{1}x^{5}$. Now we want to find the error in the $L^{\infty}$ norm of using $p_{1}$ to approximate $x\mapsto 1$. We therefore find local maxima of the error function $e_{1}(x)=1-p_{1}(x)$ on $(\ell,u)$ by setting the derivative of $e_{1}(x)$ to zero and solving for $x$. This results in the quartic equation $05b_{1}x^{4}+3b_{1}x^{2}+a=$ 0, which has closed form solutions given by the quadratic formula after the substitution $y=x^{2}$. We now let $x_{1}^{(2)}$ and $x_{2}^{(2)}$ be the solutions to this equation and let $A^{(2)}=\{\ell,x_{1}^{(2)},x_{2}^{(2)},u\}$. We repeat the procedure until $|E_{k}|:=\max\limits_{x\in[\ell,u]}|1-p_{k}(x)|\approx\max\limits_{x\in[\ell,u]}|1-p_{k+1}(x)|:=|E_{k+1}|$.
@@
We note that the matrix appearing in (18) is a Vandermonde matrix. Vandermonde matrices become notoriously ill-conditioned as the degree grows large [15, Section 4.6]. However, since in our setting we choose $d$ to be small, there is no ill-conditioning. Instead, we observe ill-conditioning when $\ell\approx u$. However, as $\ell/u\to 1$ the optimal polynomial will converge to the polynomial $\frac{x/u}{8}\left(15-10(x/u)^{2}+3(x/u)^{4}\right)$, which can be verified by noting that as $\ell/u\to 1$ all equioscillation points $x_{0},x_{1},x_{2},x_{3}$ must converge to $u$. For general $d=2q+1$, the polynomial will converge to $(x/\ell)h(1-(x/\ell)^{2})$ where $h\in\mathbb{P}_{q}$ is the $[q/0]$ Padé approximant to $(1-x)^{1/2}$ [25]. In fact, this polynomial is extremely close to the optimal polynomial for sufficiently large $\ell$. To see this, let $p^{\star}$ be the optimal approximation from $\mathbb{P}_{5}^{\operatorname*{odd}}$ and let $p(x)=\frac{x/u}{8}\left(15-10(x/u)^{2}+3(x/u)^{4}\right)$. Then,
@@
where we invoked [25, Theorem 3.1] and the fact that $p^{\star}$ is the optimal approximation to $x\mapsto 1$ from $\mathbb{P}_{5}^{\operatorname*{odd}}$. Hence, when $\ell/u\geq 1-\epsilon_{d}^{1/3}$, where $\epsilon_{\text{double}}\approx 1.1\times 10^{-16}$ is the double precision machine epsilon, then $|p^{\star}(x)-p(x)|\leq 2\epsilon_{\text{double}}$. In other words, up to double precision machine epsilon, $p^{\star}$ is equal to $p$. Therefore, whenever $\ell/u\geq 1-\epsilon_{\text{double}}^{1/3}$ the algorithm simply returns $p$ as the optimal polynomial.
The algorithm is outlined in Algorithm 3. In our experiments, we never observed Algorithm 3 taking more than five iterations to converge.
@@
input: interval $[\ell,u]$ for $\ell>0$, initial trial points $x_{1}^{(1)},x_{2}^{(1)}\in[\ell,u]$.
@@
$E_{0}=\infty$, $E_{-1}=-\infty$

### v3

@@ 5.2  Training GPT-2
In our final experiment, we compare the performance of using our Polar Express method given in Algorithm 1 inside the Muon algorithm versus Jordan’s [25] and You’s [10] methods. [^6] Our code is available at https://github.com/modichirag/GPT-opt/tree/polar, in the polar branch. We train two different GPT-2 models:
$\texttt{GPT-Small}:$ $\quad n_{\text{embd}}=768,\quad n_{\text{layer}}=12,\quad n_{\text{head}}=12$
$\texttt{GPT-Large}:$ $\quad n_{\text{embd}}=1280,\quad n_{\text{layer}}=36,\quad n_{\text{head}}=20$
and a vocabulary size of 50,25750{,}257, using a context length of $1024$. Training is performed on 1B tokens from the FineWeb dataset [2], using a batch size of 32 and a single epoch. All models are trained with mixed precision (bfloat16) on 4 H100 GPUs. For all methods we use the learning rate schedule proposed in [24], consisting of a constant phase for the first 40% of training steps followed by a linear decay. All methods for the matrix sign computations are performed in float16b precision and use five iterations.
We apply Muon selectively to certain layers of the model. Following the nano-gpt implementation [24], we assign Muon to all parameters with at least two dimensions (typically weight matrices, and excluding RMS norm parameters), excluding the embeddings, unembeddings, and the positional encodings. These excluded parameters are instead optimized with AdamW.
@@
We also experimented with adding weight decay $0.1$ to the model, keeping all else the same, in Figure 7. Here we find again that muon-PolarExp achieves a better validation and training loss for every learning rate, except one ($lr=10^{-2}$) where its performance matches that of muon-Jordan.
Acknowledgements
This work was partially supported by NSF awards 2045590 and 2234660. Computations were run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation.
@@ Acknowledgements
The aim of this section is to prove Theorem 4.1. We begin with a result that provides a few essential properties for the the polynomial solving (7) when $T=1$. This result is known as Chebyshev’s theorem [11] or the equioscillation theorem [50, Chapter 10].
@@
By Chebyshev’s Theorem [1, 11, 13] it is sufficient to show that $\mathbb{P}_{d}^{\operatorname*{odd}}$ satisfies the Haar condition: any non-zero $p\in\mathbb{P}_{d}^{\operatorname*{odd}}=\mbox{span}\{x,\ldots,x^{3},\ldots,x^{2q+1}\}$ can have at most $q$ roots in $\ell,u]$.
Since $\deg(p)=d=2q+1$ we know that $p$ can have at most $2q+1$ roots in $\mathbb{R}$. However, since $p(0)=0$ and p(x)=p(x)p(x)=-p(-x) we know that $p$ has one root at zero, and the remaining roots come in symmetric pairs (x,x)(x,-x). Because of this, $p$ can have at most $q$ roots in the positive orthant, and thus it can have at most $q$ roots in $\ell,u]\subset(0,\infty)$. Hence, $\mathbb{P}_{d}^{\operatorname*{odd}}$ satisfies the Haar condition, which yields the desired result.
@@
Note that $e(x)$ has at most $2q$ stationary points in $\mathbb{R}$, since its derivative $e^{\prime}(x)$ is a polynomial of degree $2q$. Furthermore, since $p^{\star}$ is odd, we have that $e^{\prime}(x)=-p^{\prime}(x)$ is even of degree $2q$, and thus can have at most $q$ stationary points contained in $(0,+\infty)$. Hence, there can be at most $q$ stationary points of $e(x)$ inside the interval $\ell,u]$.
By Lemma A.1 there are $q+2$ points $x_{0},\ldots,x_{q+1}\in[\ell,u]$ where $e(x)$ is maximized or minimized in $\ell,u]$. These points are either stationary points or they are endpoints of the interval $\ell,u]$. Let $n_{\text{ext}}$ be the number of stationary points and $n_{\text{stat}}$ be the number of endpoints in the set $\{x_{0},\ldots,x_{q+1}\}$. Since a point can be both a stationary point and an endpoint we have $q+2\leq n_{\text{end}}+n_{\text{stat}}$. However, $n_{\text{end}}\leq 2$ and $n_{\text{stat}}\leq q$, which follows from the previous paragraph where we showed that there are at most $q$ stationary points of $e(x)$ in $\ell,u]$. So $n_{\text{end}}+n_{\text{stat}}\leq q+2$, and consequently we must have $n_{\text{end}}=2$ and $n_{\text{stat}}=q$, as required.
Step 2: $x=\ell$ is a maximum of $e(x)$ on the interval $\ell,u]$
By Lemma A.1 and the discussion from Step 1, we know that $|e(x)|$ is maximized at $q+2$ points inside $\ell,u]$ and $q$ of these points are contained inside the open interval $(\ell,u)$. Hence, $x=\ell$ must either be a maximum or a minimum of $e(x)$. We will show that $x=\ell$ must be a maximum by contradiction.
Suppose $x=\ell$ was a minimum of $e(x)$ on $\ell,u]$. First note that $p^{\star}$ is trivially non-negative on $\ell,u]$, or else $p(x)=0$ would be a better polynomial. Hence, since $p^{\star}(0)=0$ we must have ${p^{*}}^{\prime}(\delta)>0$ for some $\delta\in[0,\ell]$, or else the zero polynomial $p(x)=0$ would be a better approximation. Hence, for some $\delta\in[0,\ell]$ we have $e^{\prime}(\delta)<0$.
We must also have $e^{\prime}(\ell)\geq 0$ or else $x=\ell$ is not a minimum of $e(x)$. Since $e^{\prime}(\delta)<0$ for some $\delta\in[0,\ell]$ and $e^{\prime}(\ell)\geq 0$, by the intermediate value theorem there exists a point $x^{*}\in[0,\ell]$ such that $e^{\prime}(x^{*})=0$. However, by the discussion above we know that all stationary points of $e$ are contained inside the open interval $(\ell,u)$. Hence, $x=\ell$ cannot be a minimum of $e(x)$ on $\ell,u]$. However, by Step 1 we know that the endpoints of $\ell,u]$ must be either minima or maxima of $e(x)$. Hence, $x=\ell$ is a maximum of $e(x)$ on $\ell,u]$.
@@
Since $e(x)$ has a maximum in $\ell,u]$ at $x=\ell$, we have $p^{\star}(\ell)=\min\limits_{x\in[\ell,u]}p^{\star}(x)$. The other two equalities are immediate consequences of the equioscillation property of $p^{\star}$ Lemma A.1 and that $x=\ell$ is a minimum of $p^{\star}$ over the set $\ell,u]$. ∎
@@
We now proceed with the proof of (10), which will be by induction. The proof for $T=1$ is an immediate consequence of Lemma A.2 and we also have $p^{\star}(\ell)=\ell_{2}$ by (11). Now suppose the result is true for all $t\leq T-1$. For $t=1,\ldots,T-1$, note that the image of $p_{t}$ on $\ell_{t},u_{t}]$ is exactly $\ell_{t+1},u_{t+1}]$ by i). Hence, if we define $g(x):=p_{T-1}\circ\cdots\circ p_{1}(x)$, then the image of $g$ on $\ell,u]$ is $\ell_{T},u_{T}]$. Furthermore, by i) we also have $g(\ell)=\ell_{T}$. Pick any $f$ such that $f\neq g$ and
@@
for some $\widetilde{p}_{1},\ldots,\widetilde{p}_{T-1}\in\mathbb{P}_{d}^{\operatorname*{odd}}$. Let the image of $f$ on $\ell,u]$ be $[a,b]$. We will prove that $\frac{a}{b}\leq\frac{\ell_{T}}{u_{T}}$ by contradiction.
Suppose $\frac{a}{b}>\frac{\ell_{T}}{u_{T}}$. Define $c=\frac{2}{a+b}$. Then, the image of the scaled function $cf$ on $\ell,u]$ is $[ca,cb]$ and $cf$ satisfies
@@
Then Algorithm 2 returns $\bm{X}_{T}=p^{\star}(\bm{M})$. Let $h\in\mathbb{P}_{q}$ be $[q/0]$ Padé-approximant to $(1-x)^{-1/2}$ [28, Section 3] and define $p(x)=xh(1-x^{2})\in\mathbb{P}_{d}^{\operatorname*{odd}}$. Define $f=p\circ\cdots\circ p$ as the composition of $p$ with itself $T$ times. Then, by Theorem 4.1, [28, Theorem 3.1], and $f(x)\geq 0$ for $x\geq 0$ we have
@@ Appendix D Remez algorithm
In this section, we show in detail how to solve (14). By Theorem 4.1, this also gives a solution to (7). We give a closed form solution for $d=3$. We then describe how the Remez algorithm [42, 43] can be used to approximate $p_{t}$ for arbitrary $d$. We finish with Algorithm 3, a simplified version of Remez for solving (14) with $d=5$. Recall (14):
@@
We begin with the case when $d=3$. In this case, there is a simple closed form for the optimal odd polynomial $p^{\star}\in\mathbb{P}_{3}^{\operatorname*{odd}}$ as described in Section 4.2. On a given interval $\ell,u]$, the optimal approximation to the constant function $x\mapsto 1$ is given by the scaled and shifted Newton-Schulz polynomial $p_{\operatorname*{NS}}(x)=\frac{3}{2}x-\frac{1}{2}x^{3}$:
@@
$p^{\star}(x)=\beta p_{\operatorname*{NS}}(\alpha x),\text{ where }\alpha=\sqrt{\frac{3}{u^{2}+\ell u+\ell^{2}}}\text{ and }\beta=\frac{4}{2+\ell u(\ell+u)\alpha^{3}}.$
One can verify that this polynomial satisfies the equioscillation condition from Lemma A.1 at $x=\ell,\frac{1}{\alpha},u$ as given in (15), with $\sqrt{-a/(3b)}=1/\alpha$ and $E=\beta-1$. Therefore, it must necessarily be the optimal approximation from $\mathbb{P}_{3}^{\operatorname*{odd}}$. Note that when $u=1$, the function $x\mapsto p_{\mathrm{NS}}(\alpha x)$ is the same polynomial derived by Chen and Chow [12].
@@
Unfortunately, for larger $d$, finding closed form expressions for optimal approximations from $\mathbb{P}_{d}^{\operatorname*{odd}}$ becomes challenging, and we know of no closed form solution. However, we can approximate the optimal polynomial using the Remez algorithm. Let $d=2q+1$. Again recalling Lemma A.1, the optimal polynomial must satisfy the equioscillation property at a set of $q+2$ points, as in (15). The Remez algorithm finds the equioscillation points $A=\{x_{0},\ldots,x_{q+1}\}$ from Lemma A.1 by iteratively refining a sequence of trial points $A^{(k)}=\{x_{0}^{(k)},\ldots,x_{q+1}^{(k)}\}$ so that $A^{(k)}$ converges to $A$. From the sequence of trial points $A^{(k)}$ the algorithm also finds a sequence of polynomials $p^{(k)}$ so that $p^{(k)}$ converges to the optimal polynomial. The convergence is very fast, and usually 10 iterations is sufficient to converge to the optimal polynomial up to double precision machine epsilon [42]. More commonly, the Remez algorithm is used to find optimal polynomial approximations to general continuous functions where $d\approx 100$ or even $d\approx 1000$. However, because the polynomial we build to approximate $\operatorname*{sign}(x)$ is a composition of polynomials, each of which has a low degree, in our setting the degree $d$ is small, usually $d=5$. For $d=5$ the Remez algorithm simplifies significantly. We now describe this simplified algorithm.
@@
We first choose an initial set of trial points $A^{(1)}$, which ideally should come close to satisfying the equioscillation property. From Lemma A.1, the unique optimal approximation $p^{\star}\in\mathbb{P}_{5}^{\operatorname*{odd}}$ satisfies the equioscillation property at four points in $\ell,u]$. Since the function we wish to approximate is constant, the equioscillation points must be extrema of $p^{\star}$ on $\ell,u]$. Because $p^{\star}$ is a odd quintic, it can have at most two local extrema on the positive real line, and thus at most two local extrema on $\ell,u]$. The other two equioscillation points must therefore be the endpoints $\ell$ and $u$. Since we know that $\ell$ and $u$ must be among the true equioscillation points, we always include them in our set of trial points. For notational simplicity, we call the other two points $q$ and $r$. We initialize $q_{1}=\frac{1}{4}\ell+\frac{3}{4}u$ and $r_{1}=\frac{3}{4}\ell+\frac{1}{4}u$, since we observe that as $\ell\to u$ these are approximately the other two equioscillation points.
We now show how to refine a candidate set of trial points $A^{(k)}=\{\ell,q_{k},r_{k},u\}$ to produce $A^{(k+1)}=\{\ell,q_{k+1},r_{k+1},u\}$ as well as an approximately equioscillating polynomial $p_{k}$. For any fixed set of trial points, we can find a degree-5 odd polynomial $p_{k}(x)=a_{k}x+b_{k}x^{3}+c_{k}x^{5}$ that satisfies
$p_{k}(\ell)=1-E_{k},\quad p_{k}(q_{k})=1+E_{k},\quad p_{k}(r_{k})=1-E_{k},\quad p_{k}(u)=1+E_{k}$ (18)
for some $E_{k}$ by solving a linear system in $a_{k},b_{k},c_{k}$ and $E_{k}$. This can be rewritten as follows:
$\begin{bmatrix}\ell&\ell^{3}&\ell^{5}&1\\ q_{k}&q_{k}^{3}&q_{k}^{5}&-1\\ r_{k}&r_{k}^{3}&r_{k}^{5}&1\\ u&u^{3}&u^{5}&-1\end{bmatrix}\begin{bmatrix}a_{k}\\ b_{k}\\ c_{k}\\ E_{k}\end{bmatrix}=\begin{bmatrix}1\\ 1\\ 1\\ 1\end{bmatrix}$. (19)
If $A^{(k)}$ were the extrema of the error function $e_{k}(x)=1-p_{k}(x)$ on $\ell,u]$, then they would be an equioscillating set for $p_{k}$, and $p_{k}$ would be the solution. Therefore, to refine $A^{(k)}$, we find the extrema of $e_{k}(x)=1-p_{k}(x)$. These can occur at $\ell,u$ and the roots of $e_{k}^{\prime}(x)$. Setting $e_{k}^{\prime}(x)=0$ yields the quartic equation $5c_{k}x^{4}+3b_{k}x^{2}+a_{k}=0$, whose two solutions are given explicitly by the quadratic formula after the substitution $y=x^{2}$. We set $q_{k+1}$ and $r_{k+1}$ to be the solutions to this equation and let $A^{(k+1)}=\{\ell,q_{k+1},r_{k+1},u\}$. We repeat the procedure until $|E_{k}|:=\max\limits_{x\in[\ell,u]}|1-p_{k}(x)|\approx\max\limits_{x\in[\ell,u]}|1-p_{k+1}(x)|=:|E_{k+1}|$.
@@
We note that the matrix appearing in (19) is a Vandermonde matrix. Vandermonde matrices become notoriously ill-conditioned as the degree grows large [17, Section 4.6]. However, since in our setting we choose $d$ to be small, there is no ill-conditioning due to large degrees. Instead, we observe ill-conditioning when $\ell\approx u$. However, as $\ell/u\to 1$ the optimal polynomial will converge to the polynomial $\frac{x/u}{8}\left(15-10(x/u)^{2}+3(x/u)^{4}\right)$, which can be verified by noting that as $\ell/u\to 1$ all equioscillation points $x_{0},x_{1},x_{2},x_{3}$ must converge to $u$. For general $d=2q+1$, the polynomial will converge to $(x/\ell)h(1-(x/\ell)^{2})$ where $h\in\mathbb{P}_{q}$ is the $[q/0]$ Padé approximant to $(1-x)^{1/2}$ [28]. In fact, this polynomial is extremely close to the optimal polynomial for sufficiently large $\ell$. To see this, let $p^{\star}$ be the optimal approximation from $\mathbb{P}_{5}^{\operatorname*{odd}}$ and let $p(x)=\frac{x/u}{8}\left(15-10(x/u)^{2}+3(x/u)^{4}\right)$. Then,
@@
where we invoked [28, Theorem 3.1] and the fact that $p^{\star}$ is the optimal approximation to $x\mapsto 1$ from $\mathbb{P}_{5}^{\operatorname*{odd}}$. Hence, when $\ell/u\geq 1-\epsilon_{d}^{1/3}$, where $\epsilon_{\text{double}}\approx 1.1\times 10^{-16}$ is the double precision machine epsilon, then $|p^{\star}(x)-p(x)|\leq 2\epsilon_{\text{double}}$. In other words, up to double precision machine epsilon, $p^{\star}$ is equal to $p$. Therefore, whenever $\ell/u\geq 1-\epsilon_{\text{double}}^{1/3}$ the algorithm simply returns the Padé approximant (that is, the scaled Newton-Schulz polynomial).
The full algorithm is given in Algorithm 3. In our experiments, we never observed Algorithm 3 taking more than five iterations to converge. This algorithm is implemented in full in Appendix G.
@@
input: interval $\ell,u]$ for $u>\ell>0$.
@@
$q_{1}=\frac{1}{4}\ell+\frac{3}{4}u,\quad r_{1}=\frac{3}{4}\ell+\frac{1}{4}u$.
$E_{0}=\infty,\quad E_{-1}=$-$\infty$